A man sits in a cannon barrel and lights the fuse. A force of 6000N propels him out of the barrel (which is 2.5m long) in 0.35s. The man has a mass of 85kg. What is the man’s velocity as he leaves the cannon?
1
Expert's answer
2019-10-18T10:09:02-0400

Let the man’s velocity as he leaves the cannon be "v", time of the force action be "t" , the force acting on the man be "F" , mass of the man be "m" .

Impulse of the man when he leaves the barrel is

"J=m\cdot v=F\cdot t," therefore

"v=\frac{F\cdot t}{m}=6000\cdot0.35/85\approx 24.7 m/s"

Answer: "24.7 m\/s."

Conditions of the problem contradict each other:

Since the force "F" is constant, the acceleration of the man should be constant and time to leave the barrel should be "t=\frac{2l}{v}=\frac{2\cdot2.5}{24.7}\approx0.2s."